What is (-5 mod 4) – (-3 mod 4) congruent to?

a) 2 mod 4

b) 3 mod 4

c) 1 mod 4

d) -1 mod 4


1
Expert's answer
2021-11-17T00:31:01-0500

It follows that

(-5 mod 4) – (-3 mod 4) = (-5 + 3) mod 4

= -2 mod 4 = (4 - 2) mod 4 = 2 mod 4


Answer: a)